You are here: Foswiki>Commissioning Web>Com20090314 (09 Oct 2009, NormCushing)Edit Attach

March BG Com Night 8, UT 2009 March 14

Observers: Dave Thompson (@LBT), Nancy Ageorges, Walter Seifert, Richard Green, Andrew Rakich (in Tucson)
Telescope Operator: Aaron Ceranski
SW Support: Tony Edgin
Telescope Support:

Summary:

  • Startup TCS problems. Lost 4.5 hours
  • Mirror cooling found to be not operating correctly.
  • Took 10 more datapoints for AGW calibration, sent numbers to Jesper.
  • 07:35UT closed dome, never reopened.

Details:

UT 02:20 Dome open. Tout=-3.9C and falling , RH=78.2% and rising, Wind~4m/s. T(M1)=0.86C. Fairly clear outside.

02:30 Tried to preset but the authorize/preset commands are not getting through...odd. Called John. Restarted IIF. The authorize complete for left LUCIFER not coming back, even after 3 minutes, but right NONE starts and completes OK. Checked that M3 is OK...it appears to be. Tony is looking into the SW.

02:47 Tony says all but ECS appear hung (no logging to syslogs). He suggests a full reset, but we suggest calling Norm just to be sure. Yes, shutdown and restart all subsystems. Tony found 4-5 PMCRs, 2 ECSs, very strange. Traced to too rapid a network restart (netconfig) on the servers.

03:19 Restarting the subsystems. All but OSS and PSFR had trouble, both trying to start on 105. Two instances of IRS found on 105, killing them.

03:25 Authorize still not getting through. Subsystems do not appear to be logging to their respective syslogs. We tried sending the telescope to a different position with the MCS GUI (MCS is on 107) and it looks like it is working normally, with the telescope moving and MSSPU updating appropriately.

03:34 Tony is calling Chris. An authorize request from an xterm is returning:
LBTO@lbtmu105:15 % irc Authorize 'bentGregorianFront left' LUCIFER
irc - InfraRed Camera Client V.1.3.1
Connection refused either by IIF or IRS. Both services must be running in order to use irc.
IRS must be running...(re)starting IRS on 105 (service irs start). ICE exception encountered.
03:45 Calling Stephen.

03:54 Changed verbose mode in irtc_client.config from 2 to 0 (needed for proper iraf script operation).

04:09 authorize from xterm command line returns "Exit 66" multiple times.

04:22 authorize from xterm command line returns "Exit 162".

04:25 Calling Jose. John arrives.

04:33 We note the left M1 is cooling slowly at best (-4C outside, M1 is still only 0.5). John is looking into it. Chillers working, but no glycol seems to be getting to the mirror cooling circuit. JRousseau awakened to check with Aaron. Pressure blower found to be not running. Manual power switch, breakers checked. Internal reset button did the trick.

05:23 Preset from LUCIFER gets through...

05:35 The pointing offset appears to be back! But ends up being 170" in the other direction. Dave Ashby arrives back from the mountain and tells us that the SX elevation encoder tape has been damaged beyond repair or use, and could account for some issues.

In the end, Jose cannot restart IRS on lbtmu105, so IRAF scripts won't play. For now, LUCIFER will execute the presets.

~07:00 Beginning data acquisition for AGw offset measurements. Using field of BS9143, starting at PA=90. Jesper connects from Potsdam. Seeing settles down to ~0.7" on guide camera, with rms wfe ~ 200 nm or better.

#ID  UTTIME   PCS_X_____PCS_Y ACQ  LUCI Xcenter Ycenter
###============================================================
 14 06:39:45  108.261  18.399  71  0001 1049.91 1030.40  Preset a while before 
 14 06:54:08  108.262  18.416  72  0002 1049.55 1030.38  fresh preset
 22 06:57:29  102.075  98.413  73  0003 1064.08 1022.70
 25 07:01:07   88.105 143.999  74  0004 1069.59 1016.98
 24 07:06:01  -97.489 130.471  75  0005 1038.69  977.58
 17 07:08:58 -171.959  33.233  76  0006 1013.97  975.38
 13 07:12:14  -32.015  15.976  77  0007 1032.61  998.14
 21 07:16:39   82.290  93.498  78  0008 1058.84 1016.33
 26 07:21:00   30.433 173.592  79  0009 1063.62  995.34
 19 07:24:10   -0.336  68.876  80  0010 1052.14 1005.20 
 18 07:28:29 -113.529  39.387  81  0011 1023.56  982.71 

07:35 Closing dome for approaching radar echo.

10:43 Still cloudy. Tout=-5.0, RH=84%, Wind~4m/s. T(M1)=-2.25

12:18 Clear, but now >90% humidity. Temp still falling and humidity still rising, so Tucson folks called it a night.

Performance verification

TCS SW tests needing to be performed

BG Berformance Verification


Plan (copied from prior nights):

Re-Measure AGW-PCS Transformation (2hr per field; Storm, Thompson):

  • Prerequisites:
    LUCIFER accurately oriented on sky - yes (rms now 0.029 degrees)
  • Use source list: LBTdefault.cat (it has been modified to include this field)
  • Pointto M67astrom at PA=0 or BS9177 at PA=??
  • New finding chart for M67astrom here:
    http://wiki.lbto.arizona.edu/twiki/bin/view/Commissioning/StarCharts
  • Measure:
    1. Take LUCIFER rotator center data (30s ea, TRACK mode, PA=0, 180, 90, 270) (done on 20090311 UT)
    2. Observe multiple (~20) guide stars at PA=0, some repeats
    3. Repeat field at PA=15, significant fraction of stars from previous step
    4. Do two stars at PA=70, 50, 30, 10, -10, -30, -50, -70, different radii
    5. Repeat LUCIFER rotator center data (30s ea, TRACK mode, PA=0, 180, 90, 270)
  • Record:
    1. Date and time of observation
    2. 30s LUCIFER image, filename
      (may need J or narrow band for bright stars, coadds)
    3. Star number observed, acquisition image for this preset
    4. PCS XY coordinates
    5. AGW XY and encoder values

Field Aberration Measurements

Persson's Star = Start time =

Guide Star x y PA Doug's Angle wfscimage000*
- 0 612.5 0 9 1-3
6 0 520.0 43 273 4-6
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        

End time =

Persson's Star =
Start time =

Guide Star x y PA Doug's Angle wfscimage000*
- 0 612.5 0 9 1-3
6 0 520.0 43 273 4-6
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        

End time =

Pupil Measurements

Please check that the optics alignment remains OK with the pupil viewer in LUCIFER.

  • Alignment of chief ray to LUCIFER pupil with M3 (Andrew, Walter, John, 2 hours)
    • Point to bright star
    • Insert LUCIFER pupil viewer
    • Take Pupil image
    • Offset rotator 180 deg
    • Take Pupil image
    • Iterate M3 tip-tilt to minimize the pupil motion
    • Compensate pointing with IE,CA

Measure LUCIFER full-field images above and below focus.

-- JoarBrynnel - 02 Mar 2009
Topic revision: r18 - 09 Oct 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback